Five $20 bills, four $10 bills, four $5 bills and seven $1 bills are placed in a bag. If a bill is chosen at random, what is the expected value for the amount chosen?
Answer #1

Ans Number 1 Bello ! $20 $10 $ 5 $! 20 L Expected value of amount= ax 20)+ ( 4 x 10 = 5+2+1+= 8 +0.35=8.35 Expected value of